OpenStudy (anonymous):

A garden has sqrt 13 width and length 7 sqrt 13. What is the perimeter of the garden in simplest radical form?

OpenStudy (anonymous):

P = 2 ( L + w ) = 2 ( √13 + 7√13 ) = 16√13